### v2

# Non-Heisenbergian Quantum Mechanics

Abstract
Relaxing the postulates of an axiomatic theory is a natural way to find more general theories, and historically, the discovery of non-Euclidean geometry is a famous example of this procedure. Here, we use this way to extend quantum mechanics by ignoring the heart of Heisenberg’s quantum mechanics—We do not assume the existence of a position operator that satisfies the Heisenberg commutation relation, $[\hat{x},\hat{p}]=i\hbar$. The remaining axioms of quantum theory, besides Galilean symmetry, lead to a more general quantum theory with a free parameter $l_{0}$ of length dimension, such that as $l_{0}\to 0$ the theory reduces to standard quantum theory. Perhaps surprisingly, this non-Heisenbergian quantum theory, without a priori assumption of the non-commutation relation, leads to a modified Heisenberg uncertainty relation, $\Delta x\Delta p\geq\sqrt{\hbar^{2}/4+l_{0}^{2}(\Delta p)^{2}}$, which ensures the existence of a minimal position uncertainty, $l_{0}$, as expected from various quantum gravity studies. By comparing the results of this framework with some observed data, which includes the first longitudinal normal modes of the bar gravitational wave detector AURIGA and the $2S-1S$ transition in the hydrogen atom, we obtain upper bounds on the $l_{0}$.
I Introduction
There are various motivations to establish a theory of quantum mechanics without a position operator. Firstly, as shown by Pauli, there is no self-adjoint time operator conjugate to Hamiltonian operator Pauli (1958); Muga et al. (2007); Egusquiza and Muga (1999), and therefore, if one considers a self-adjoint position operator conjugate to momentum operator, space and time would not be on equal footing—a fact that is not in harmony with special relativity Toller (1999); Terno (2014); Caban et al. (2014), see also  [^1] In fact, there is no position operator that satisfies following natural demands Newton and Wigner (1949); Caban et al. (2014); Terno (2014): i) Its components be self-adjoint and commute with each other, ii) it satisfies canonical commutation relation with momentum operator, iii) It transforms like a three-vector under rotations, and iv) It transforms consistently under Lorentz boosts. The operator satisfying all of the conditions but the last condition is obtained by Newton and Wigner Newton and Wigner (1949): An eigen-state of Newton-Wigner position operator in one reference frame, which represents a localized state at a point, is not localized in other reference frames. Hence, the Newton-Wigner position operator have no simple covariant meaning Newton and Wigner (1949); Barr et al. (2023).. Secondly, and more importantly, as shown by Hegerfeldt, strictly localized states lead to faster-than-light probability propagation Hegerfeldt (1974); Hegerfeldt and Ruijsenaars (1980); Hegerfeldt (1985). This means that, in a self-consistent relativistic quantum mechanics, exactly localized states could not exist Thaller and Thaller (1984); Barat and Kimball (2003). Therefore, the position operator is not a well-defined concept if the eigenstates of the position operator represent exactly localized states. Although exactly localized states are not possible, unsharp localization is possible Bracken and Melloy (1999); Krekora et al. (2004, 2004); Kazemi et al. (2018), which can be described using a proper positive-operator-valued measure instead of a self-adjoint position operator Terno (2014); Kazemi et al. (2018). In fact, even in the non-relativistic setting, recent progress in the Wigner-Araki-Yanase theorem shows that exactly localized states are not practically accessible via position measurements Kuramochi and Tajima (2023); Busch and Loveridge (2011). Another motivation for avoiding exactly localized states comes from quantum gravity. Although a complete and agreed-upon theory of quantum gravity is still lacking, various approaches to quantum gravity Hossenfelder (2013) as well as gedankenexperiments that heuristically combine quantum theory and general relativity Calmet et al. (2004); Das and Vagenas (2008); Scardigli (1999); Adler and Santiago (1999); Amelino-Camelia suggest the existence of a limit to observable spatial resolution.
The aim of the present work is to show how one can establish such a theory of quantum mechanics without a position operator. Although we avoid the position operator, (at least unsharp) position measurements are routinely performed in quantum experiments, even in the relativistic regime Sinha et al. (2010); Pleinert et al. (2021); Korzh et al. (2020). Therefore, a complete quantum theory must be able to predict the observed position distributions. In fact, a space-time discription of quamtum phenomena is critical for understanding the relativistic aspects of quantum information processes Caban et al. (2014); Peres and Terno (2004) and addressing some fundamental question in quantum theory Krekora et al. (2004); Wagner et al. (2011, 2012); Glasgow and Ware (2020); e.g., does photons satisfy uncertainty relations? If yes, how it could be proven without a reference to a position operator? Terno (2014); Bialynicki-Birula and Bialynicka-Birula (2012). Our starting point is the modern axiomatic formulation of (generalized) quantum theory in which the notion of observable is not limited to self-adjoint operators; an observable-distribution is given by a positive-operator-valued measure (POVM), not necessary a projection-valued measure (PVM) corresponding to a self-adjoint operator Masanes et al. (2019); Busch et al. (2009); Holevo (2011); Busch et al. (1989); Kunjwal et al. (2014). A justification for this generalization is provided by Gleason’s theorem Busch et al. (1996); Busch (2003), which guarantees that this structure is the most general one compatible with the probabilistic interpretation of quantum mechanics (for more details, see Busch et al. (1996)). Another justification, based on an operational approach to quantum theory, can be found in Masanes et al. (2019). In fact, the use of POVMs has been previously suggested to describe various observable distributions which their description via self-adjoint operators is problematic. Examples include the arrival time distribution Giannitrapani (1997); Hegerfeldt et al. (2010); Hegerfeldt and Muga (2010), work distribution Roncaglia et al. (2014), screen observable Werner (1986), and so on Busch et al. (2009). The aim of the present work is to establish a space-resolved quantum mechanics in which the position probability distribution is given by a proper position POVM at the fundamental level. Although the use of POVMs to describe particle localization has been previously suggested Busch et al. (1989), as far as we know, this idea has not been fully extended to systematically develop a complete quantum theory—including the derivation of its consequences on uncertainty relations, time evolution, the spectrum of interacting particle systems, and so on. Note that, in conventional quantum mechanics, the position operator not only determines the position probability density but also serves other roles, especially in canonical quantization where it is used to determine the Hamiltonian operator corresponding to a classical Hamiltonian. We will discuss how a proper position POVM could be consistently used for this purpose as well.
In this work, our focus is on the non-relativistic regime; however, the results shed light on the subject in the general case. The relativistic generalization will be reported in future works. In this regard, the rest of the paper is organized as follows. In Section II, we obtain the general position probability density consistent with Galilean symmetry and the general structure of quantum theory. In Section III, we demonstrate that this general position probability density leads to a generalized position-momentum uncertainty relation, as well as a minimum position uncertainty, denoted as $l_{0}$. In Section IV, we discuss the time evolution in the presence of external potential and construct the associated Schrödinger equation without a position operator. Specifically, we study the simple harmonic oscillator in this framework and obtain an upper bound on $l_{0}$ by comparing the theoretical results with experimental data of the bar gravitational wave detector AURIGA. In Section V, we systematically generalize this formalism to interacting multi-particle systems, and compute the spectrum of the hydrogen atom in the presence of minimal position uncertainty. Using experimental data on $1S-2S$ transition, we derive a more stringent upper bound on $l_{0}$. In Section VI, we summarize the results and provide a discussion on future prospects.
II Probability density and Galilean symmetry
According to generalized quantum mechanics Busch et al. (1989), we assume that, for any $\bm{x}\in\mathbb{R}^{3}$, there is an operator $\hat{Q}_{\bm{x}}$ such that the position probability density at $\bm{x}$ is given by
$\rho_{t}(\bm{x})=\langle\psi_{t}|\hat{Q}_{\bm{x}}|\psi_{t}\rangle.$ (1)
In the conventional Heisenberg quantum mechanics, $\hat{Q}_{\bm{x}}$ is a projection operator constructed out of the eigenstates of a position operator, as $|\bm{x}\rangle\langle\bm{x}|$. We do not consider this assumption. Based on the probability theory, it is not necessary that $\hat{Q}_{\bm{x}}$ be a projection operator, it is only needed that $\hat{Q}_{\bm{x}}$ be a positive operator, $\hat{Q}_{\bm{x}}\geq 0$, and the set of $\hat{Q}_{\bm{x}}$ be normalized, i.e. $\int\hat{Q}_{\bm{x}}d^{3}x=\mathbb{1}$. These conditions make $\{\hat{Q}_{\bm{x}}|\bm{x}\!\in\!\mathbb{R}^{3}\}$ a POVM.
Now, let’s find the general position probability density for a free spinless particle consistent with non-relativistic space-time symmetries, including spatial and temporal translations, rotation, and Galilean boost. The space and time translation symmetries lead to $\hat{Q}^{t}_{\bm{x+a}}=e^{-i\bm{\hat{P}}.\bm{a}/\hbar}\hat{Q}^{t}_{\bm{x}}e^{i\bm{\hat{P}}.\bm{a}/\hbar}$, and $\hat{Q}^{t+\tau}_{\bm{x}}=e^{i\hat{H}\tau}\hat{Q}^{t}_{\bm{x}}e^{-i\hat{H}\tau/\hbar},$ respectively, which for infinitesimal transformations reduce to
$[\hat{Q}^{t}_{\bm{x}},\hat{\bm{P}}]=-i\hbar\nabla\hat{Q}^{t}_{\bm{x}},$ (2)
and
$[\hat{Q}^{t}_{\bm{x}},\hat{H}]=+i\hbar\partial_{t}\hat{Q}^{t}_{\bm{x}},$ (3)
where $\hat{\bm{P}}/\hbar$ and $\hat{H}/\hbar$ are generators of space and time translation, respectively, and $\hat{Q}^{t}_{\bm{x}}$ is time-dependent position POVM in the Heisenberg picture, $\hat{Q}^{t}_{\bm{x}}\equiv e^{i\hat{H}t/\hbar}\hat{Q}_{\bm{x}}e^{-i\hat{H}t/\hbar}$. The equation (2) could be considered as an alternative for the usual position-momentum commutation relation. Using space and time translations symmetries, the $\hat{Q}^{t}_{\bm{x}}$ could be written in momentum representation as:
$\hat{Q}^{t}_{\bm{x}}\!=\!\!\int\!\!\ Q(\bm{p},\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}\left[e^{i\hat{H}t/\hbar}|\bm{k}\rangle\langle\bm{p}|e^{-i\hat{H}t/\hbar}\right]d^{3}\bm{p}d^{3}\bm{k},$
where, $Q(\bm{p},\bm{k})=\langle\bm{k}|\hat{Q}^{0}_{\bm{0}}|\bm{p}\rangle$. Similarly, the Galilean boost and rotation symmetries lead to the following commutation relations, respectively:
$[\hat{Q}^{t}_{\bm{x}},\hat{\bm{K}}]=-i\hbar\ t\nabla\hat{Q}^{t}_{\bm{x}},$ (4)
$[\hat{Q}^{t}_{\bm{x}},\hat{\bm{J}}]=-i\hbar\ \bm{x}\!\times\!\nabla\hat{Q}^{t}_{\bm{x}}.$ (5)
Here, $\hat{\bm{K}}/\hbar$ and $\hat{\bm{J}}/\hbar$ are generators of the boost and rotation transformations in Hilbert space. Then, using the Lie algebra of the projective representation of the Galilean group Weinberg (1995); Ballentine (2014), the equations (5) and (4) lead to the following constraints on the $Q(\bm{p},\bm{k})$ function:
$(\nabla_{\bm{p}}$ $+$ $\nabla_{\bm{k}})Q(\bm{p},\bm{k})=0,$
$(\bm{p}\times\nabla_{\bm{p}}$ $+$ $\bm{k}\times\nabla_{\bm{k}})Q(\bm{p},\bm{k})=0,$
which together ensure the $Q(\bm{p},\bm{k})$ is a function of $|\bm{p}-\bm{k}|$. Therefore, the general bilinear position probability density as (1) and consistent with non-relativistic space-time symmetries, Eq. (2-5), is given by
$\rho_{t}(\bm{x})\!=\!\!\!\!\int\!f(|\bm{p}-\bm{k}|)\tilde{\psi}_{t}(\bm{p})\tilde{\psi}_{t}^{*}(\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$ (6)
where $\tilde{\psi}_{t}(\bm{p})$ is momentum space wave function, $\langle\bm{p}|\psi_{t}\rangle$, and $f(|\bm{p}-\bm{k}|)$ is an isotropic positive-definite kernel. The kernel $f(|\bm{p}-\bm{k}|)$ is also called a radial positive definite function. The characterization of radial positive definite functions is a classical result of I. Schoenberg Schoenberg (1938): A function $f(r)$ is radial positive definite on $\mathbb{R}^{n}$, If and only If there exists a finite positive measure $\alpha$ on $\mathbb{R}^{+}$ such that
[…]
where $\Omega_{n}(r)=\Gamma(n/2)(2/r)^{(n-2)/2}J_{(n-2)/2}(r),$ and $J_{n}$ is the Bessel function of order $n$ (see Wells and Williams (2012); Gneiting (1999); Wendland (2004) for further details). It is easy to see that the probability density which is given in (6) is conserved; it satisfies continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ where the probability current $\bm{J}$ is given by
$\bm{J}(\bm{x})=\int\frac{\bm{p}+\bm{k}}{2m}f(|\bm{p}-\bm{k}|)\tilde{\psi}(\bm{p})\tilde{\psi}^{*}(\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$ (7)
for a free particle with $\hat{H}=\hat{\bm{P}}^{2}/2m$. Note that the only difference between the equations (6) and (7) and their counterparts in conventional quantum theory is the existence of the kernel $f(|\bm{p}-\bm{k}|)$.
III Modified uncertainty relation
To facilitate a detailed comparison with conventional quantum mechanics, let us compute the moments of the position probability distribution, $\langle x_{i}^{n}\rangle=\int x_{i}^{n}\rho(\bm{x})d^{3}\bm{x}$. Using the identity $\int e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}x=(2\pi\hbar)^{3}\delta(\bm{p}-\bm{k}),$ and eq.(6) we find that,
$\langle x_{i}^{n}\rangle=(2\pi\hbar)^{3}\sum_{j=0}^{n}\frac{n!(i\hbar)^{j}}{(n-j)!j!}\xi_{j}\int x_{i}^{(n-j)}|\psi(\bm{x})|^{2}d^{3}x,$
where $\xi_{j}\equiv\frac{\partial^{j}}{\partial p^{j}_{i}}f(|\bm{p}-\bm{k}|)|_{\bm{p}=\bm{k}}$, and $\psi(\bm{x})$ is defined using momentum wave function as $\psi(\bm{x})\equiv(2\pi\hbar)^{-3/2}\int\tilde{\psi}(\bm{p})e^{i\bm{x}.\bm{p}/\hbar}d^{3}p$, as usual. Note that, the $\psi(\bm{x})$ should not be interpreted as a ”wave function” in the usual sense, because without a position operator and position eigenstates, $\langle\bm{x}|\psi\rangle$ cannot be defined. Moreover, in the general case, the position probability density is given by eq. (6), not $|\psi(\bm{x})|^{2}$.
The normalization condition, $\int\rho(\bm{x})d^{3}x=1$, ensures that $\xi_{0}=(2\pi\hbar)^{-3}$. Moreover, using the fact that $f$ is a positive kernel, it is easy to prove that $f(x)\leq f(0)$ Wells and Williams (2012), which ensures that $\xi_{1}=0$ and $\xi_{2}\leq 0$. Therefore, if $l_{0}$ is defined as $l_{0}=\sqrt{-\xi_{2}(2\pi\hbar)^{3}\hbar^{2}}$, it is easy to see that,
$\langle x_{i}\rangle$ $=$ $\int x_{i}|\psi(\bm{x})|^{2}d^{3}\bm{x},$
$\langle x^{2}_{i}\rangle$ $=$ $l_{0}^{2}+\int x^{2}_{i}|\psi(\bm{x})|^{2}d^{3}\bm{x},$
where $l_{0}$ is a positive constant with length dimension. Therefore, it is clear that,
$(\Delta x_{i})^{2}=l_{0}^{2}+\int(x_{i}-\langle x_{i}\rangle)^{2}|\psi(\bm{x})|^{2}d^{3}x,$ (8)
in which $\Delta x_{i}=\sqrt{\langle x_{i}^{2}\rangle-\langle x_{i}\rangle^{2}}$. Note that, according to the fact that the quantity $\int(x_{i}-\langle x_{i}\rangle)^{2}|\psi(\bm{x})|^{2}d^{3}x$ can be arbitrarily small, the equation (8) ensures that $(\Delta x_{i})_{\text{Inf}}=l_{0}$, where $(\Delta x_{i})_{\text{Inf}}=\text{Inf}\{\Delta^{\psi}x_{i}|\ |\psi\rangle\in\mathcal{H}\}$.
It should be noted that the assumption that an arbitrary sharp localized state exists, $l_{0}=0$, determines the kernels $f$ uniquely as a constant function $f(|\bm{p}-\bm{k}|)=f(0)=(2\pi\hbar)^{-3}$ Kijowski (1974), which means that the equation (6) reduces to $\rho=|\psi|^{2}$. Moreover, if $f=f(0)$, the equation (7) reduces to the usual Schrödinger probability current, i.e. $\bm{J}=\frac{\hbar}{m}\Im[\psi^{*}\nabla\psi]$. In fact, in this case, the current formalism reduces to the standard Heisenberg quantum theory.
The existence of a minimum position uncertainty modifies the usual position-momentum uncertainty relation. Using properties of Fourier transformation, we know that
$\int(x_{i}-\langle x_{i}\rangle)^{2}|\psi(\bm{x})|^{2}d^{3}x\int(p_{i}-\langle p_{i}\rangle)^{2}|\tilde{\psi}(\bm{p})|^{2}d^{3}p\geq\hbar^{2}/4,$
which, together with (8), leads to the following modified position-momentum uncertainty relation:
$\Delta x_{i}\Delta p_{i}\geq\sqrt{\hbar^{2}/4+l_{0}^{2}(\Delta p_{i})^{2}}.$ (9)
This modified uncertainty relation is the same as the common generalized uncertainty relation at the first order, $\Delta x_{i}\Delta p_{i}\geq\hbar/2(1+2l_{0}^{2}(\Delta p_{i})^{2}/\hbar^{2})$, which appears in a wide range of approaches to quantum gravity Amati et al. (1989); Maggiore (1993a); Scardigli (1999); Das and Vagenas (2008); Hossenfelder (2013). Often, such modification is seen as a manifestation of a deformed canonical commutator Maggiore (1993b); Kempf et al. (1995); Fadel and Maggiore (2022) which has been related to a deformation of space-time symmetry Fadel and Maggiore (2022); Maggiore (1994); Bosso et al. (2023a). However, it is important to emphasize that we did not assume any modified symmetry or deformed commutation relation. The general form of quantum probability density (6) and the modified uncertainty relation (9) are derived from the Galilean symmetry.
In the quantum gravity literature, it is often assumed that the value of $l_{0}$ is near the Planck scale ($l_{p}=10^{-35}$ m). Nonetheless, it is worth mentioning that some theories with varying gravitational coupling, such as those with additional dimensions Arkani-Hamed et al. (1998); Antoniadis et al. (1998); Hossenfelder et al. (2003), imply a larger effective Planck length. In our framework, $l_{0}$ is a free parameter, and we do not presuppose any specific value for it. In principle, the value of $l_{0}$ should be determined from a direct comparison of the theoretical result with experimental data Pikovski et al. (2012).
IV Single particle in an external potential
To find an experimental bound on $l_{0}$, we need to extend the model to non-free particles. Firstly, we consider a single particle in the presence of an external potential and then the model is extended for an interacting many-particle system. For a particle in an external potential, we assume that the position probability density is given by a POVM as $\rho_{t}(\bm{x})=\langle\psi_{t}|\hat{Q}_{\bm{x}}|\psi_{t}\rangle$, and the time evolution of the state is given by the following Schrödinger equation:
$i\hbar\frac{d}{dt}|\psi\rangle=\hat{H}|\psi\rangle,$ (10)
where $\hat{H}=\hat{\bm{P}}^{2}/2m+\hat{V}$ and the potential operator, associated with the classical scalar potential $V(\bm{x})$, is naturally defined as follows:
$\hat{V}=\int V(\bm{x})\hat{Q}_{\bm{x}}d^{3}x.$ (11)
Note that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator, i.e., $\hat{V}=\int V(\bm{x})|\bm{x}\rangle\langle\bm{x}|d^{3}x$. In the general case, $l_{0}\neq 0$, this definition is conceptually consistent with our definition of position probability density, eq.(6), due to the fact that this definition ensures
$\langle\psi|\hat{V}|\psi\rangle=\int\rho^{\psi}(\bm{x})V(\bm{x})d^{3}x.$ (12)
In fact, the definition (11), on the other hand, can be deduced by assuming Eq.(12).
It is easy to show that, the same as in the free particle case, the Galilean symmetry implies that $\langle\bm{p}|\hat{Q}_{\bm{x}}|\bm{k}\rangle=f(|\bm{p}-\bm{k}|)e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}.$ Therefore, in the momentum representation, the modified Schrödinger equation (10) reads
$i\hbar\frac{\partial}{\partial t}\tilde{\psi}(\bm{p})=\frac{\bm{p}^{2}}{2m}\tilde{\psi}(\bm{p})+\int\tilde{U}(\bm{p}-\bar{\bm{p}})\tilde{\psi}(\bar{\bm{p}})d^{3}\bar{\bm{p}},$ (13)
where $\tilde{U}(\bm{p}-\bar{\bm{p}})=f(|\bm{p}-\bar{\bm{p}}|)\int V(\bm{x})e^{i(\bm{p}-\bar{\bm{p}}).\bm{x}/\hbar}.$ The equations (6) and (13) together ensure the local probability conservation—they lead to the continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ in which the probability current $\bm{J}$ is given by
$\bm{J}(\bm{x})=\int\left(\alpha(\bm{p}+\bm{k})+\beta(\bm{p}-\bm{k})\right)e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$
where $\alpha=f(|\bm{p}-\bm{k}|)\tilde{\psi}^{*}(\bm{k})\tilde{\psi}(\bm{p})/2m$, and
$\beta=\frac{f(|\bm{p}-\bm{k}|)}{(\bm{p}-\bm{k})^{2}}\times$
$\left(\!\tilde{\psi}^{*}(\bm{k})\!\!\int\!\!\tilde{U}(\bm{p}-\bar{\bm{p}})\tilde{\psi}(\bar{\bm{p}})d^{3}\bar{\bm{p}}\!-\!\tilde{\psi}(\bm{p})\!\!\int\!\!\tilde{U}^{*}\!(\bm{k}-\bar{\bm{k}})\tilde{\psi}^{*}(\bar{\bm{k}})d^{3}\bar{\bm{k}}\right)\!\!.$
It is easy to check that the given probability current is correctly transformed under Galilean transformations Brown and Holland (1999). In principle, this probability current may be experimentally probed using the measurement of arrival time distribution Ali et al. (2003, 2007); Vona et al. (2013); Tumulka (2022); Roncallo et al. (2023); Ayatollah Rafsanjani et al. (2023); Kazemi and Hosseinzadeh (2023).
The only difference between the Eq.(13) and the usual Schrödinger equation in momentum representation is the existence of the factor $f$, which can be effectively considered as deformation of the potential. However it is important to remark that, the physical content of this equation is different from the usual Schrödinger equation with a deformed potential, due to the presence of the factor $f$ in the definition of the position probability density, Eq.(6). Nonetheless, the equations (6) and (13) together lead to
$m\frac{d^{2}\langle\bm{x}\rangle}{dt^{2}}=-\langle\nabla V\rangle,$ (14)
where $\langle\nabla V\rangle$ is naturally defined as $\int\nabla V(\bm{x})\rho(\bm{x})d^{3}x$—which is exactly same as the result in standard quantum theory.
For a more detailed comparison with standard quantum mechanics, it is useful to study the stationary states of the harmonic oscillator. For a d-dimensional harmonic oscillator with $V(\bm{x})=\frac{m\omega^{2}}{2}\sum_{i}^{d}x_{i}^{2}$, the Hamiltonian eigenvalue problem $\hat{H}|\psi\rangle=E|\psi\rangle$, for any $f(|\bm{p}-\bm{k}|)$ kernel, leads to
$E\tilde{\psi}(\bm{p})=\frac{\bm{p}^{2}}{2m}\tilde{\psi}(\bm{p})-\frac{m\omega^{2}}{2}\hbar^{2}\nabla^{2}_{\bm{p}}\tilde{\psi}(\bm{p})+\frac{d}{2}ml_{0}^{2}\omega^{2}\tilde{\psi}(\bm{p}),$
which is the same as the standard equation except for the last term. Therefore, in the momentum representation, the harmonic stationary states are the same as standard ones, while the eigen-energies are given by
$E_{\bm{n}}=\sum_{i}^{d}\left(n_{i}\!+\!\frac{1}{2}\right)\hbar\omega+\frac{d}{2}ml_{0}^{2}\omega^{2}.$ (15)
Here $n=(n_{1},..,n_{d})$ and $n_{i}$ are integer numbers. Note that, although the energy eigenstates are the same as energy eigenstates in usual quantum theory and lead to the same momentum distribution, the position probability density is different from the usual, due to the presence of $f$ kernel in Eq.(6). Moreover, from Eq.(14), it is clear that the time evolution of the position expectation value $\langle x\rangle$ is the same as in standard quantum theory, and is consistent with classical motion in the classical limit—the fact that is in sharp contrast with the results of deformed-algebra formalisms, which are used to experimental probing of the deformed commutation relation Bawaj et al. (2015); Kumar and Plenio (2020). Nonetheless, in our model, the higher-order moments of the position probability density $\langle x^{m}\rangle$ and their time evolution are different from usual cases. To probe these effects some clever experimental design is needed: For example, see the optomechanical setups that are proposed to probe the effect of minimal position uncertainty near the Planck scale Pikovski et al. (2012); Kumar and Plenio (2018); Chevalier et al. (2022). Detailed study of such setups in our non-Heisenbergian quantum theory is beyond the scope of the present work and are left for future studies. Nonetheless, as a first estimation, using Eq.s (15), it is essay to show that
$m\omega^{2}l_{0}^{2}+\hbar\omega/2\leq\langle p^{2}\rangle/2m+m\omega^{2}\langle x^{2}\rangle/2,$ (16)
which could be used to estimate an upper limit on $l_{0}$  [^2] Note that, although the potential (and thus the Hamiltonian) is defined up to a constant $C$, the inequality (16) is independent of this constant. If one uses a constantly shifted Hamiltonian, $\hat{H}^{\prime}=\hat{H}+C$, although the Hamiltonian eigenvalues are shifted, the inequality (16) remains correct.. In Refs. Marin et al. (2013, 2014), an experimental limit for the modal energy, i.e., the right-hand side of (16), has been reported by analyzing the residual motion of the first longitudinal mode of the AURIGA gravitational wave bar detector, which is cooled down to sub-millikelvin Vinante et al. (2008). The measured modal energy, $E_{\text{exp}}\!=\!1.3\!\times\!10^{26}$, leads to an upper limit for the minimal position uncertainty as $l_{0}<3.8\times 10^{16}l_{P}$ Marin et al. (2013). Nevertheless, it should be noted that the direct application of the one-particle theory to the center of mass of a macroscopic object with a very large mass, such as the AURIGA detector with $m=10^{13}M_{p}$, may be questionable—at least, the value of $l_{0}$ may not be same for all objects Bosso et al. (2023b); Chevalier et al. (2022). In principle, for a more accurate consideration, a multi-particle extension of our framework is needed. In the next section, we will develop such an extension, and based on it, we obtain a more justifiable upper limit on $l_{0}$ using the observed Hydrogen atom spectrum.
V Interacting particles system
In this section, we extend the theory to multi-particle systems. For simplicity, we only consider a two-particle system; however, extending the results to an $N$-particle system is straightforward. The Hilbert space of the system is considered, as usual, as the tensor product of one-particle Hilbert spaces. The probability density of a two-particle system in configuration space is given by
$\rho_{t}(\bm{x}_{1},\bm{x}_{2})=\langle\psi|\hat{Q}^{t}_{\bm{x}_{1},\bm{x}_{2}}|\psi\rangle,$ (17)
where, $\{\hat{Q}^{t}_{\bm{x}_{1},\bm{x}_{2}}|(\bm{x}_{1},\bm{x}_{2})\in\mathbb{R}^{6}\}$ is a POVM in the two-particle Hilbert space. Similar to the single-particle theory, the non-relativistic space-time symmetries lead to the following commutation relations:
$\left[\hat{Q}^{t}_{\bm{x}},\hat{H}\right]$ $=$ $+i\hbar\partial_{t}\hat{Q}^{t}_{\bm{x}},$ (18)
$\left[\hat{Q}^{t}_{\bm{x}},\hat{\bm{P}}\right]$ $=$ $-i\hbar\sum_{i=1}^{N}\nabla_{\bm{x}_{i}}\hat{Q}^{t}_{\bm{x}},$ (19)
$\left[\hat{Q}^{t}_{\bm{x}},\hat{\bm{K}}\right]$ $=$ $-i\hbar\ t\sum_{i=1}^{N}\nabla_{\bm{x}_{i}}\hat{Q}^{t}_{\bm{x}},$ (20)
$\left[\hat{Q}^{t}_{\bm{x}},\hat{\bm{J}}\right]$ $=$ $-i\hbar\sum_{i=1}^{N}\bm{x}_{i}\!\times\!\nabla_{\bm{x}_{i}}\hat{Q}^{t}_{\bm{x}},$ (21)
where, $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$. It is easy to check that the above commutation relations lead to $\langle\bm{k}_{1},\bm{k}_{2}|\hat{Q_{\bm{0}}^{0}}|\bm{p}_{1},\bm{p}_{2}\rangle=f(|\bm{p}_{1}-\bm{k}_{1}|,|\bm{p}_{2}-\bm{k}_{2}|)$, where $f$ is a positive semi-definite kernel in $\mathbb{R}^{6}$. If, as in usual quantum theory, we assume that the separable multi-particle states lead to a separable two-particle probability density—i.e., $|\psi\rangle=\prod_{i}|\psi_{i}\rangle\Rightarrow\rho(\bm{x})=\prod_{i}\rho_{i}(\bm{x}_{i})$—then, we have
$f(|\bm{p}_{1}-\bm{k}_{1}|,|\bm{p}_{2}-\bm{k}_{2}|)=f_{1}(|\bm{p}_{1}-\bm{k}_{1}|)f_{2}(|\bm{p}_{2}-\bm{k}_{2}|),$
where $f_{1}$ and $f_{2}$ are radial-positive-kernels in $\mathbb{R}^{3}$. Therefore, the two-particle probability density in configuration space is given by
$\rho_{t}(\bm{x})\!=\!\!\!\!\int\!\prod_{i}f_{i}(|\bm{p}_{i}-\bm{k}_{i}|)\tilde{\psi}_{t}(\bm{p})\tilde{\psi}_{t}^{*}(\bm{k})e^{i(\bm{p}_{i}-\bm{k}_{i}).\bm{x}_{i}/\hbar}d^{3}\bm{p}_{i}d^{3}\bm{k}_{i},$
where $\bm{x}=(\bm{x}_{1},\bm{x}_{2})$, $\bm{p}=(\bm{p}_{1},\bm{p}_{2})$, and $\bm{k}=(\bm{k}_{1},\bm{k}_{2})$. Note that we consider a system made of two different particles, which, in principle, could have different quantum behaviors characterized by different kernels.
We assume that the time evolution of the state is given by the Schrödinger equation, $i\hbar\frac{d}{dt}|\psi\rangle=\hat{H}|\psi\rangle$, in which $\hat{H}=\sum_{i=1}^{2}\frac{\hat{\bm{P}}_{i}^{2}}{2m_{i}}+\hat{V}$ and the potential operator $\hat{V}$, associated with the classical potential $V(\bm{x}_{1},\bm{x}_{2})$, is defined as follows:
$\hat{V}=\int V(\bm{x}_{1},\bm{x}_{2})\hat{Q}_{\bm{x}_{1},\bm{x}_{2}}d^{3}\bm{x}_{1}d^{3}\bm{x}_{2}.$
Note that, when $\hat{Q}_{\bm{x}}$ are projection operators, this definition reduces to the standard potential operator—i.e., $\hat{V}=\int V({\bm{x}_{1},\bm{x}_{2}})|{\bm{x}_{1},\bm{x}_{2}}\rangle\langle{\bm{x}_{1},\bm{x}_{2}}|d^{3}x_{1}d^{3}x_{2}$. The Galilean symmetry limits the possible forms of the interaction potential—from $[\hat{\bm{K}},\hat{H}]=i\hbar\hat{\bm{P}}$ and Eq.(20), or from $[\hat{\bm{P}},\hat{H}]=0$ and Eq.(19), it is easy to see that,
$(\nabla_{\bm{x}_{1}}+\nabla_{\bm{x}_{2}})V(\bm{x}_{1},\bm{x}_{2})=0.$ (22)
Moreover, $[\hat{\bm{J}},\hat{H}]=0$ together with Eq.(21) lead to
$(\bm{x}_{1}\times\nabla_{\bm{x}_{1}}+\bm{x}_{2}\times\nabla_{\bm{x}_{2}})V(\bm{x}_{1},\bm{x}_{2})=0.$ (23)
The equations (22) and (23) imply that the potential must be a function of $|\bm{x}_{2}-\bm{x}_{1}|$, which is the same condition that arises from Galilean symmetry in conventional quantum mechanics Weinberg (2015).
As a crucial example, we investigate the energy levels of the hydrogen atom within the introduced framework. The hydrogen atom may be considered as a two-particle system with interaction potential $V(|\bm{x}_{1}-\bm{x}_{2}|)=-e^{2}/4\pi\epsilon_{0}|\bm{x}_{1}-\bm{x}_{2}|$. In momentum representation, the energy eigenvalue problem $E|\Psi\rangle=\hat{H}|\Psi\rangle$ can be written as:
$E\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})$ $=$ $\left(\frac{\bm{p}^{2}_{1}}{2m_{1}}+\frac{\bm{p}^{2}_{2}}{2m_{2}}\right)\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})$
$+$ $\int\tilde{U}(\bm{p}_{1}-\bm{k}_{1},\bm{p}_{2}-\bm{k}_{2})\tilde{\Psi}(\bm{k}_{1},\bm{k}_{2})d^{3}\bm{k}_{1}d^{3}\bm{k}_{2},$
where, $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\langle\bm{p}_{1},\bm{p}_{2}|\Psi\rangle$, and
$\tilde{U}\!\!=\!\!f_{1}(|\bm{p}_{1}|)f_{2}(|\bm{p}_{2}|)\!\!\!\int\!\!V(|\bm{x}_{1}-\bm{x}_{2}|)e^{\frac{i}{\hbar}(\bm{p}_{1}.\bm{x}_{1}+\bm{p}_{2}.\bm{x}_{2})}d^{3}\bm{x}_{1}d^{3}\bm{x}_{2}.$
It is well-known that, in the case of two interacting particles, working with the total momentum $\bm{p}_{c}=\bm{p}_{1}+\bm{p}_{2}$ and the relative momentum $\bm{p}_{r}=(m_{1}\bm{p}_{2}-m_{2}\bm{p}_{1})/(m_{1}+m_{2})$ is useful. After separation of the external and internal degrees of freedom, $\tilde{\Psi}(\bm{p}_{1},\bm{p}_{2})=\tilde{\psi}_{c}(\bm{p}_{c})\tilde{\psi}_{r}(\bm{p}_{r})$, the two-body problem can be split into $E_{c}\psi_{c}(\bm{p}_{c})=\frac{\bm{p}^{2}_{c}}{2M}\psi_{c}(\bm{p}_{c})$, and
$E_{r}\tilde{\psi}_{r}(\bm{p}_{r})=\frac{\bm{p}^{2}_{r}}{2\mu}\tilde{\psi}_{r}(\bm{p}_{r})+\int\tilde{U}_{r}(\bm{p}_{r}-\bm{k}_{r})\tilde{\psi}_{r}(\bm{k}_{r})d^{3}\bm{k}_{r},$ (25)
where $\tilde{U}_{r}(\bm{p})=(2\pi\hbar)^{3}f_{1}(|\bm{p}|)f_{2}(|\bm{p}|)\int V(|\bm{x}|)e^{i\bm{p}.\bm{r}/\hbar}d^{3}\bm{r}$, $E=E_{c}+E_{r}$, $M=m_{1}+m_{2}$, $\mu=m_{1}m_{2}/(m_{1}+m_{2})$, and $\bm{r}$ represent the relative coordinate, $\bm{x}_{2}-\bm{x}_{1}$. Using the Taylor expansion of $f_{i}$, in the first-order approximation, Eq.(25) reduces to
$E\psi_{r}(\bm{r})\!=\!-\frac{\hbar^{2}}{2\mu}\!\nabla^{2}\psi_{r}(\bm{r})\!-(1+\frac{l_{1}^{2}+l_{2}^{2}}{2}\nabla^{2})V(r)\psi_{r}(\bm{r}),$
where $\psi_{r}(\bm{r})=(2\pi\hbar)^{-3/2}\int\tilde{\psi}_{r}(\bm{p})e^{i\bm{p}.\bm{r}/\hbar}d^{3}p$, and $l_{1}$ and $l_{2}$ are minimal position uncertainty of the particles, which are given by $l_{i}^{2}=(2\pi\hbar)^{3}\hbar^{2}f^{\prime\prime}_{i}(0)$. Applying the usual perturbation theory, the leading corrections to the $S$-energy levels are given by
$\Delta E_{n}=-\frac{e^{2}}{2a_{0}n^{2}}\frac{16\pi(l_{1}^{2}+l_{2}^{2})}{na_{0}^{2}},$
in which $a_{0}=4\pi\epsilon_{0}\hbar^{2}/\mu e^{2}$ is the Bohr radius. Comparing this result with the high-precision experimental data of hydrogen atom spectroscopy can lead to an upper bound for the minimal position uncertainties. To achieve this, we focus on the extremely narrow $1S-2S$ two-photon transition. In recent years, the frequency of the $1S-2S$ transition has been accurately measured using a cesium clock with a relative uncertainty of $4.5\times 10^{-15}$ Matveev et al. (2013); Parthey et al. (2011). This relative uncertainty leads to the following upper bound on the minimal position uncertainties:
$l_{i}\leq\sqrt{l_{1}^{2}+l_{2}^{2}}\leq 2.8\times 10^{16}l_{p},$
which is in agreement with the upper bound obtained in the previous section using AURIGA data.
Finally, it is important to remark that the corrected energy levels in our framework differ from the ones in the deformed commutation relation models Stetsko and Tkachuk (2006); Bouaziz and Ferkous (2010); Quesne and Tkachuk (2010). Especially, in contrast to our result, in some deformed commutation relation models, the hydrogen energy spectrum depends on the center-of-mass momentum Quesne and Tkachuk (2010). In principle, these differences can be used to experimentally distinguish between these models and our non-Heisenbergian quantum mechanics.
VI Summary and out-looks
Understanding the core content of quantum mechanics requires us to disentangle the hidden logical relationships between the basics assumptions and results of this theory. One of the most important assumptions of Heisenberg quantum mechanics is that there exists a position self-adjoint operator that satisfies the canonical commutation relation. This assumption is usually considered necessary to derive most of the important results of the theory, especially the uncertainty relation. Nonetheless, in this work, by ignoring the position operator, we demonstrate that Heisenberg quantum mechanics could be seen as a special case of a more general self-consistent quantum theory, which has most of fundamental characteristics of the Heisenberg quantum mechanics, such as the Galilean symmetry, probability conservation, position-momentum uncertainty relation, the discrete and bounded from bellow energy spectrum in bound states, and so on. This approach opens an important door to generalize the quantum theory consistent with a given space-time symmetry. It is important to remark that the deformation of commutation relation, as used in GUP literature Maggiore (1993b); Kempf et al. (1995); Fadel and Maggiore (2022), is not consistent with usual space-time symmetries Fadel and Maggiore (2022); Maggiore (1994); Bosso et al. (2023a)—the usual canonical commutation relation is a direct result of the Galilean and also Poincare algebras if we consider a self adjoint position operator that correctly transform under special translation. Therefore, it seems that for constructing a non-relativistic quantum theory with a non-zero minimal position uncertainty, the presented formalism is more acceptable than deformed commutation relation approach.
One of the most important extensions of the present work is its relativistic extension. It could provide a way to circumvent the fundamental problems in relativistic quantum mechanics, which arise from the impossibility of defining a completely proper relativistic position operator. In fact, due to the problems mentioned in the Introduction and some other issues Sebens (2019); Kazemi et al. (2018); Terno (2014), despite much effort in recent years Kowalski and Rembieliński (2011); Wagner et al. (2011, 2012); Bialynicki-Birula and Bialynicka-Birula (2012); Babaei and Mostafazadeh (2017); Kazemi et al. (2018); Kiessling and Tahvildar-Zadeh (2018); Sebens (2019); Hawton (2019, 2021); Glasgow and Ware (2020), a complete and agreed-upon formulation of relativistic quantum mechanics with a clear space-time resolution has not yet been proposed. Nonetheless, in relativistic quantum experiments, the position measurements are routinely performed with high accuracy (just for example see Sinha et al. (2010); Pleinert et al. (2021); Korzh et al. (2020)), and thus, a complete quantum theory must be able to predict the observed position distributions Sebens (2019); Krekora et al. (2004). We believe that a proper relativistic extension of our formulation of quantum theory without a position operator could circumvent the mentioned problems, as has been previously shown by one of the authors in a very special case Kazemi et al. (2018). The detailed study of the relativistic extension is left for future works.
Finally, it should be noted that the generalization of quantum theory without a position operator is natural and interesting from the perspective of some non-standard interpretations of quantum theory, such as Bohmian mechanics Barnea et al. (2018), Nelson stochastic mechanics Nelson (1966), and the many interacting worlds model Hall et al. (2014). In these theories, the operators are not fundamental; they are secondarily mathematical objects that are useful for computing statistical observable results Dürr et al. (2004a). However, in the formalism of these theories, the position probability density and probability current have fundamental important roles. So, the approach of the present work could shed some new light on the generalization of these alternative formulations of quantum theory, especially in the relativistic regime, which is an interesting open problem Foo et al. (2022); Dürr et al. (2014); Gisin (2018); Holland (1995); Dürr et al. (2004b).
VII Acknowledgment
We are grateful to B. Taghavi and B. Askari for useful comments.